a. Expand (π‘Ž+𝑏)5. Hence find the coefficient of π‘₯ in the expansion of (4π‘₯+2/9π‘₯)5

b. The coefficient of π‘₯2 in the expansion of (1+π‘₯)n is 45. Given that 𝑛 is a positive integer, find the value of 𝑛.


1
Expert's answer
2022-05-31T13:45:16-0400

a.


"(a+b)^5=\dbinom{5}{0}a^5+\dbinom{5}{1}a^4b+\dbinom{5}{2}a^3b^2"

"+\dbinom{5}{3}a^2b^3+\dbinom{5}{4}ab^4++\dbinom{5}{5}b^5"

"=a^5+5a^4b+10a^3b^2+10a^2b^3+5ab^4+b^5"



"(4x+\dfrac{2}{9x})^5=1024x^5+\dfrac{5(256)(2)x^4}{9x}"

"+\dfrac{10(64)(4)x^3}{81x^2}+\dfrac{10(16)(8)x^2}{729x^3}+\dfrac{5(4)(16)x}{6561x^4}"

"+\dfrac{32x^4}{59049x^5}=1024x^5+\dfrac{2560x^3}{9}+\dfrac{2560x}{81}"

"+\dfrac{1280}{729x}+\dfrac{320}{6561x^3}+\dfrac{32}{59049x^5}"

b.


"(1+x)^n=\dbinom{n}{0}+\dbinom{n}{1}x+\dbinom{n}{2}x^2+..."

"\dbinom{n}{2}=45"

"\dfrac{n(n-1)}{2}=45"

"n(n-1)=90, n>0"

Then "n=10"